# A Set of Tutorials for the LAMMPS Simulation Package [Article v1.0]

### v1

@@ 3.4.1 System preparation
These lines are used to define the most basic parameters, including the atom, bond, and angle styles, as well as interaction potential. Here, lj/cut/tip4p/long imposes a Lennard-Jones potential with a cut-off at $12\,\text{$\text{\AA{}}$}$ and a long-range Coulomb potential.
So far, the commands are relatively similar to those in the previous tutorial, Polymer in water, with two major differences: the use of lj/cut/tip4p/long instead of lj/cut/coul/long, and pppm/tip4p instead of pppm. When using lj/cut/tip4p/long and pppm/tip4p, the interactions resemble the conventional Lennard-Jones and Coulomb interactions, except that they are specifically designed for the four-point water model. As a result, LAMMPS automatically creates a four-point water molecule, assigning type O atoms as oxygen and type H atoms as hydrogen. The fourth massless atom (M) of the TIP4P water molecule does not have to be defined explicitly, and the value of $0.1546\,\text{$\text{\AA{}}$}$ corresponds to the O-M distance of the TIP4P-2005 water model [13]. All other atoms in the simulation are treated as usual, with long-range Coulomb interactions. Another novelty, here, is the use of kspace_modify slab 3.0 that is combined with the non-periodic boundaries along the $z$ coordinate: boundary p p f. With the slab option, the system is treated as periodical along $z$, but with an empty volume inserted between the periodic images of the slab, and the interactions along $z$ effectively turned off.
@@
The lattice command defines the unit cell. Here, the face-centered cubic (fcc) lattice with a scale factor of 4.04 has been chosen for the future positioning of the atoms of the walls. The region command defines a geometric region of space. By choosing $\text{xlo}=-3$ and $\text{xlo}=3$, and because we have previously chosen a lattice with a scale factor of 4.04, the region box extends from $-12.12\leavevmode\nobreak\ \text{\AA{}}$ to $12.12\leavevmode\nobreak\ \text{\AA{}}$ along the $x$ direction. The create_box command creates a simulation box with 5 types of atoms: the oxygen and hydrogen of the water molecules, the two ions ($\text{Na}^{+}$, $\text{Cl}^{-}$), and the atoms from the walls. The simulation contains 1 type of bond and 1 type of angle (both required by the water molecules). The parameters for these bond and angle constraints will be given later. The extra/ (…) keywords are for memory allocation. Finally, the labelmap commands assign alphanumeric type labels to each numeric atom type, bond type, and angle type.
@@
Within the last three lines, a region named rliquid is created based on the last defined lattice, fcc 4.04. rliquid will be used for depositing the water molecules. The molecule command opens up the molecule template called water.mol, and names the associated molecule h2omol. The new molecules are placed on the fcc 4.04 lattice by the create_atoms command. The first parameter is 0, meaning that the atom IDs from the water.mol file will be used. The number 482793 is a seed that is required by LAMMPS, it can be any positive integer.
@@
[…]
@@
The bond_coeff command, used here for the O-H bond of the water molecule, sets both the spring constant of the harmonic potential and the equilibrium bond distance of $0.9572\leavevmode\nobreak\ \text{\AA{}}$. The constant can be 0 for a rigid water molecule because the SHAKE algorithm will maintain the rigid structure of the water molecule (see below) [46, 47]. Similarly, the angle_coeff command for the H-O-H angle of the water molecule sets the force constant of the angular harmonic potential to 0 and the equilibrium angle to $104.52^{\circ}$.
Alongside parameters.inc, the groups.inc file contains several group commands to selects atoms based on their types:
@@
The groups.inc file also defines the walltop and wallbot groups, which contain the WALL atoms located in the $z>0$ and $z<0$ regions, respectively::
@@
The run 0 command runs the simulation for 0 steps, which is sufficient for creating the system and saving its state. The write_data command generates a file called system.data containing the information required to restart the simulation from the final configuration produced by this input file. With the nocoeff option, the parameters from the force field are not included in the .data file. Run the create.lmp file using LAMMPS, and a file named create.data will be created alongside create.lmp.
@@
Let us move the atoms and place them in more energetically favorable positions before starting the actual molecular dynamics simulation.
Open the equilibrate.lmp file that was downloaded alongside create.lmp during the tutorial setup. It contains the following lines:
@@
Now, let us use the SHAKE algorithm to maintain the shape of the water molecules [46, 47].
@@
Here the SHAKE algorithm applies to the O-H bond and the H-O-H angle of the water molecules. The kbond keyword specifies the force constant that will be used to apply a restraint force when used during minimization. This last keyword is important here, because the spring constants of the rigid water molecules were set to 0 (see the parameter.inc file).
@@
Let us perform an energy minization by adding the following lines to equilibrate.lmp:
@@
Let us equilibrate further the entire system by letting both fluid and piston relax at ambient temperature. Here, the commands are written within the same equilibrate.lmp file, right after the reset_timestep command.
Let us update the positions of all the atoms and use a Nosé-Hoover thermostat. Add the following lines to equilibrate.lmp:
@@
Figure 25: a) Pressure, $p$, of the nanosheared electrolyte system simulated in Tutorial 4 as a function of the time, $t$. b) Distance between the walls, $\Delta z$, as a function of $t$.
@@ 3.4.2 Imposed shearing
From the equilibrated configuration, let us impose a lateral motion on the two walls and shear the electrolyte. Open the last input file named shearing.lmp. It starts with the following lines:
@@
The setforce commands cancel the forces on walltop and wallbot. As a result, the atoms in these two groups will not experience any forces from the rest of the system. Consequently, in the absence of external forces, these atoms will conserve the initial velocities imposed by the two velocity commands.
@@
Let us also extract the density and velocity profiles using the chunk/atom and ave/chunk commands. These commands are used to divide the system into bins and return the desired quantities, here the velocity along $x$ (vx) within the bins. Add the following lines to shearing.lmp:
@@
Run the simulation using LAMMPS. The averaged velocity profile for the fluid is plotted in Fig. 26. As expected for such Couette flow geometry, the fluid velocity increases linearly along $z$, and is equal to the walls velocities at the fluid-solid interfaces (no-slip boundary conditions).
@@
[…]
Note 3.20.
The viscosity calculated at such a high shear rate may differ from the expected bulk value. In general, it is recommended to use a lower value for the shear rate. Note that for lower shear rates, the ratio of noise-to-signal is larger, and longer simulations are needed. Another important point to consider is that the viscosity of a fluid next to a solid surface is typically larger than in bulk due to interaction with the walls [50]. Therefore, one expects the present simulation to yield a viscosity that is slightly higher than what would be measured in the absence of walls.
@@ 3.5.1 Prepare and relax
The initial topology given by silica.data is a small amorphous silica structure. This structure was created using a force field called Vashishta [52]. If you open the silica.data file, you will find in the Atoms section that all silicon atoms have a charge of $q=1.1\,\text{e}$, and all oxygen atoms have a charge of $q=-0.55\,\text{e}$.
Note 3.21.
Assigning the same charge to all atoms of the same type is common with many force fields, including the force fields used in the previous tutorials. This changes once ReaxFF is used: the charge of each atom will adjust to its local environment.
@@
pair_coeff * * reaxCHOFe.inc Si O
@@
In this case, the pair_style reaxff is used without a control file. The safezone and mincap keywords are added to prevent allocation issues, which sometimes can trigger segmentation faults and bondchk errors. The pair_coeff command uses the reaxCHOFe.inc file, which should have been downloaded during the tutorial set up. Finally, the fix qeq/reaxff is used to perform charge equilibration [53], which occurs at every step. The values 0.0 and 10.0 represent the low and the high cutoffs, respectively, and $1.0\text{e}-6$ is the tolerance. The maxiter sets an upper limit to the number of attempts to equilibrate the charge.
@@
To print the averaged charges qSi and qO using the thermo_style command, and create images of the system. Add the following lines to relax.lmp:
@@
We can also use the fix reaxff/species to evaluate what species are present within the simulation. It will be useful later when the system is deformed, and bonds are broken:
@@
write_data relax.data
Run the relax.lmp file using LAMMPS. As seen from relax.species, only one species is detected, called O384Si192, representing the entire system.
@@ 3.5.2 Deform the structure
[…]
@@
pair_coeff * * reaxCHOFe.inc Si O
@@
Here, no barostat is used because the change in the box volume will be imposed by the fix deform.
@@
write_data deform.data
@@ 3.5.3 Decorate the surface
pair_coeff * * reaxCHOFe.inc Si O H
@@
Here, the $+1\mathrm{e}{-10}$ was added to the denominator of the variable qH to avoid dividing by 0 at the beginning of the simulation. Finally, let us create a loop with 10 steps, and create two hydrogen atoms at random locations at every step:
@@ 3.6 Tutorial 6: Water adsorption in silica
The objective of this tutorial is to combine molecular dynamics and grand canonical Monte Carlo simulations to compute the adsorption of water molecules in cracked silica material (Fig. 34). This tutorial illustrates the use of the grand canonical ensemble in molecular simulation, an open ensemble where the number of atoms or molecules in the simulation box can vary. By employing the grand canonical ensemble, we will set the chemical potential of water within a nanoporous SiO2 structure.
@@ 3.6.1 Generation of the silica block
The create_atoms commands are used to place 240 Si atoms, and 480 atoms, respectively. This corresponds to an initial density of approximately $2$ g/cm3, which is close to the expected final density of amorphous silica at 300 K.
Now, specify the pair coefficients by indicating that the first atom type is Si and the second is O:
@@
[…]
@@
In the third step, the system is equilibrated at the final desired conditions, $T=300\,\text{K}$ and $p=1\,\text{atm}$, using an anisotropic pressure coupling:
@@
Here, an anisotropic barostat is used. Anisotropic barostats adjust the dimensions independently, which is generally suitable for a solid phase.
Run the simulation using LAMMPS. From the «Charts» window, the temperature evolution can be observed, showing that it closely follows the desired annealing procedure (Fig. 36 a). The evolution of the box dimensions over time confirms that the box deformed during the last stage of the simulation (Fig. 36 b). After the simulation completes, the final LAMMPS topology file called generate.data will be located next to generate.lmp (Fig. 35).

### v2

@@ 3.4.1 System preparation
These lines are used to define the most basic parameters, including the atom style, the forms of the non-bonded, bond, and angle potentials, as well as other specifics of the non-bonded interactions. Here, lj/cut/tip4p/long imposes a Lennard-Jones potential with a cut-off at $12\,\text{\AA {}}$ and a long-range Coulomb potential. The parameters O, H, O-H, and H-O-H correspond respectively to the oxygens, hydrogens, O-H bonds, and H-O-H angle constraints of the water molecules; their definitions, provided by the labelmap commands, will be clarified below.
So far, the commands are relatively similar to those in the previous tutorial, Polymer in water, with two major differences: the use of lj/cut/tip4p/long instead of lj/cut/coul/long, and pppm/tip4p instead of pppm. When using lj/cut/tip4p/long and pppm/tip4p, the interactions resemble the conventional Lennard-Jones and Coulomb interactions, except that they are specifically designed for the four-point water model. As a result, LAMMPS automatically adds the fourth point to the water molecules, assigning type O atoms as oxygen and type H atoms as hydrogen. The fourth massless atom (M) of the TIP4P water molecule does not have to be defined explicitly, and the value of $0.1546\,\text{\AA {}}$ corresponds to the O-M distance of the TIP4P-2005 water model [13]. All other atoms in the simulation are treated as usual, with long-range Coulomb interactions. Another novelty, here, is the use of kspace_modify slab 3.0 that is combined with the non-periodic boundaries along the $z$ coordinate: boundary p p f. With the slab option, the system is treated as periodical along $z$, but with an empty volume inserted between the periodic images of the slab, and the interactions along $z$ effectively turned off.
@@
The lattice command defines the unit cell. Here, the face-centered cubic (fcc) lattice with a scale factor of 4.04 has been chosen for the future positioning of the atoms of the walls. The region command defines a geometric region of space. By choosing $\text{xlo}=-3$ and $\text{xlo}=3$, and because we have previously chosen a lattice with a scale factor of 4.04, the region box extends from $12.12\leavevmode\nobreak\ \text{\AA {}}$ to $12.12\leavevmode\nobreak\ \text{\AA {}}$ along the $x$ direction. The create_box command creates a simulation box with 5 types of atoms: the oxygen and hydrogen of the water molecules, the two ions ($\text{Na}^{+}$, $\text{Cl}^{-}$), and the atoms from the walls. The simulation contains 1 type of bond and 1 type of angle (both required by the water molecules). The parameters for these bond and angle constraints will be given later. The extra/ (…) keywords are for memory allocation. Finally, the labelmap commands assign alphanumeric type labels to each numeric atom type, bond type, and angle type, concepts already introduced in previous tutorials.
@@
Within the last three lines, a region named rliquid is created based on the last defined lattice, fcc 4.04. rliquid will be used for introducing the water molecules. The molecule command opens up the molecule template called water.mol, and names the associated molecule h2omol. The new molecules are placed on the fcc 4.04 lattice by the create_atoms command. The first parameter is 0, meaning that the atom IDs from the water.mol file will be used. The number 482793 is a seed that is required by LAMMPS, it can be any positive integer.
@@
Each pair_coeff assigns the depth of the LJ potential (in kcal/mol), and the distance (in Ångströms) at which the particle-particle potential energy is 0. As noted in previous tutorials, with the important exception of pair_coeff O WALL, pairwise interactions were only assigned between atoms of identical types. By default, LAMMPS calculates the pair coefficients for the interactions between atoms of different types (i and j) by using geometric average: $\epsilon_{ij}=\sqrt{\epsilon_{ii}\epsilon_{jj}}$, $\sigma_{ij}=\sqrt{\sigma_{ii}\sigma_{jj}}$. However, if the default value of $1.472\,\text{kcal/mol}$ was used for $\epsilon_{\text{O-WALL}}$, the solid walls would be extremely hydrophilic, causing the water molecules to form dense layers. As a comparison, the water-water energy $\epsilon_{\text{O-O}}$ is only $0.185199\,\text{kcal/mol}$. Therefore, to make the walls less hydrophilic, the value of $\epsilon_{\text{O-WALL}}$ was reduced.
@@
The bond_coeff command, used here for the O-H bond of the water molecule, sets both the spring constant of the harmonic potential and the equilibrium bond distance of $0.9572\leavevmode\nobreak\ \text{\AA {}}$. The force constant can be 0 for a rigid water molecule because the SHAKE algorithm, which will be used in the input at a later step, will constrain the intramolecular structure of the water molecules (see below) [47, 48]. Similarly, the angle_coeff command for the H-O-H angle of the water molecule sets the force constant of the angular harmonic potential to 0 and the equilibrium angle to $104.52^{\circ}$.
Alongside parameters.inc, the groups.inc file contains several group commands to define groups of atoms based on their types:
@@
The groups.inc file also defines the walltop and wallbot groups, which contain the WALL atoms located in the $z>0$ and $z<0$ regions, respectively:
@@
The run 0 command initializes the simulation but does not advance positions or velocities which is required for cleanly saving the state. The write_data command generates a file called system.data containing the information required to restart the simulation from the final configuration produced by this input file. With the nocoeff option, the parameters from the force field are not included in the .data file. Run the create.lmp file using LAMMPS, and a file named create.data will be created alongside create.lmp.
@@
Let us move the atoms and place them in more energetically favorable positions before starting the actual molecular dynamics simulation. Open the equilibrate.lmp file that was downloaded alongside create.lmp during the tutorial setup. Same as before, it contains the following lines:
@@
Now, let us use the SHAKE algorithm to maintain the shape of the water molecules [47, 48] by adding the following line to the script.
@@
Here the SHAKE algorithm applies to the O-H bond and the H-O-H angle of the water molecules. The kbond keyword specifies the force constant that will be used to apply a restraint force when used during minimization. This last keyword is important here, because the spring constants of the rigid water molecules were set to 0 (see the parameters.inc file).
Note 3.33.
LAMMPS provides several ways to keep molecules rigid during a simulation. The fix shake command is appropriate and efficient for constraining individual bonds or bonds and angles within small molecules like water while using a per-atom time integration fix command like fix nve or fix nvt. However, it fails for linear molecules like CO2 or constraining larger or more complex objects. In such cases, the fix rigid family of commands can be used to perform time integration for translation and rotation of groups of atoms as rigid bodies.
@@
Let us perform an energy minimization by adding the following lines to equilibrate.lmp:
@@
Let us equilibrate further the entire system by letting both fluid and wall relax at ambient temperature. Here, the commands are written within the same equilibrate.lmp file, right after the reset_timestep command.
Let us do a molecular dynamics simulation using the Nosé-Hoover thermostat. Add the following lines to equilibrate.lmp:
@@
Note 3.34.
Just like the undump command can cancel an active dump, other objects defined in a LAMMPS input script can be cancelled when no longer needed. For example, you can use unfix to remove a previously defined fix, and uncompute to delete a compute.
@@
Figure 25: a) Pressure, $p$, of the nanosheared electrolyte system simulated in Tutorial 4 as a function of the time, $t$. b) Distance between the walls, $\Delta z$, as a function of $t$. The orange line shows the raw data, and the blue line represents a time-averaged curve.
@@ 3.4.2 Imposed shearing
From the equilibrated configuration, let us impose a lateral motion on the two walls and shear the electrolyte. Open the last input file named shearing.lmp. It starts as follows:
@@
The setforce commands cancel the forces on walltop and wallbot in the $x$ direction. As a result, the atoms in these two groups will not experience any forces along $x$ from the rest of the system. Consequently, in the absence of external forces, these atoms will conserve the initial velocities imposed by the two velocity commands. As seen previously, although the forces on these atoms are set to zero, the fix setforce still stores the forces acting on the group before cancellation, which can later be extracted for analysis (see below).
@@
Let us also extract the density and velocity profiles using the chunk/atom and ave/chunk commands. When deployed as below, these commands discretize the simulation domain into spatial bins and compute and output average properties of the atoms belonging to each bin, here the velocity along $x$ (vx) within the bins. Add the following lines to shearing.lmp:
@@
Run the simulation using LAMMPS. The averaged velocity profile for the fluid is plotted in Fig. 26. As expected for such a Couette flow geometry, the fluid velocity increases linearly along $z$, and is equal to the walls velocities at the fluid-solid interfaces (no-slip boundary conditions).
@@
where $\tau$ is the stress applied by the fluid on the shearing wall, and $\dot{\gamma}$ the shear rate [50]. Here, the shear rate is approximately $\dot{\gamma}=20\cdot 10^{9}\,\text{s}^{-1}$ (Fig. 26), the average force on each wall is given by f_mysf1[1] and f_mysf2[1] and is approximately $2.7\,\mathrm{kcal/mol/\AA {}}$. Using a surface area for the walls of $A=6\cdot 10^{-18}\,\text{m}^{2}$, one obtains an estimate for the shear viscosity for the confined fluid of $\eta=3.1\,\mathrm{mPa\cdot s}$ using Eq. (2).
Note 3.36.
The viscosity calculated at such a high shear rate may differ from the expected bulk value. In general, it is recommended to use a lower value for the shear rate. Note that for lower shear rates, the signal-to-noise ratio is smaller, and longer simulations are needed. Another point to consider is that the viscosity of a fluid next to a solid surface is typically larger than in bulk due to interaction with the walls [51]. Therefore, one expects the present simulation to yield a viscosity that is slightly higher than what would be measured in the absence of walls.
@@ 3.5.1 Prepare and relax
The initial topology given by silica.data corresponds to a small amorphous silica structure. This structure was generated in a prior simulation using the Vashishta force field [53]. If you open the silica.data file, you will find in the Atoms section that all silicon atoms have a charge of $q=1.1\,\text{e}$, and all oxygen atoms have a charge of $q=-0.55\,\text{e}$.
Note 3.37.
Assigning the same charge to all atoms of the same type is common with many force fields, including the force fields used in the previous tutorials. This changes once ReaxFF is used: the charge of each atom will adjust to its local environment through charge equilibration.
Figure 28: A slice of the amorphous silica simulated during Tutorial 5, where atoms are colored by their charges. Dangling oxygen groups appear in greenish, bulk Si atoms with a charge of about $1.8\leavevmode\nobreak\ \text{e}$ appear in red/orange, and bulk O atoms with a charge of about $0.9\leavevmode\nobreak\ \text{e}$ appear in blue.
@@
pair_coeff * * ffield.reax.CHOFe Si O
@@
In this case, the pair_style reaxff is used without a control file (see note below). The safezone and mincap keywords are added to prevent allocation issues, which sometimes can trigger segmentation faults and bondchk errors. The pair_coeff command uses the ffield.reax.CHOFe file, which should have been downloaded during the tutorial set up. Finally, the fix qeq/reaxff is used to perform charge equilibration [54], which occurs at every step. The values 0.0 and 10.0 represent the low and the high cutoffs, respectively, and $1.0\text{e}-6$ is the tolerance, i.e., the precision to which the atomic charges are equilibrated during the charge equilibration process. The maxiter sets an upper limit to the number of attempts to equilibrate the charge.
Note 3.38.
The pair_style reaxff command optionally accepts a control file, which defines control variables such as global parameters of the ReaxFF potential, as well as performance and output settings. If no control file is provided, as in this tutorial, LAMMPS uses its default values, which correspond to those in Adri van Duin’s original stand-alone ReaxFF code [14].
@@
The definition of the equal style variables qSi and qO make use of functions pre-defined within LAMMPS that allow calculating the total charge of atoms belonging to a group (charge()) and the total number of atoms in the group (count()). To print the averaged charges qSi and qO using the thermo_style command, and create images of the system. Add the following lines to relax.lmp:
@@
The fix ave/histo command samples values over a group of atoms and builds a histogram over a specified range divided into bins. In this tutorial, it is used to monitor the charge distributions of silicon and oxygen atoms. The parameters 10 500 5000 specify how often the histogram is updated and averaged, -1.5 2.5 set the value range, 1000 is the number of bins, and v_vq is the variable being histogrammed.
We can also use fix reaxff/species to evaluate what species are present within the simulation. It will be useful later when the system is deformed, and bonds are broken:
@@
write_data relax.data nofix
The write_data command is used with the nofix keyword to print a data file without extra sections from the reaxff/species command. Run the relax.lmp file using LAMMPS. As seen from relax.species, only one species is detected, called O384Si192, representing the entire system.
Note 3.39.
With the aniso keyword, the three dimensions of the simulation box can change independently. This is particularly relevant for solids and other systems where anisotropic stresses may develop.
@@ 3.5.2 Deform the structure
Let us apply a deformation to the structure to force some $\text{Si}-\text{O}$ bonds to break (and eventually re-assemble). Open the deform.lmp file, which must contain the following lines:
@@
pair_coeff * * ffield.reax.CHOFe Si O
@@
Here, no barostat is used because the change in the box volume will be imposed by the fix deform, see below.
@@
write_data deform.data nofix
The fix deform command applies a continuous deformation by elongating the simulation box along the x-axis at a constant engineering shear strain rate, specified by erate, of $5\times 10^{-5}\leavevmode\nobreak\ \text{fs}^{-1}$.
@@
Figure 32: Amorphous silicon oxide after deformation during Tutorial 5. The atoms are colored by their charges. Dangling oxygen groups appear in greenish, bulk Si atoms with a charge of about $1.8\leavevmode\nobreak\ \text{e}$ appear in red/orange, and bulk O atoms with a charge of about $0.9\leavevmode\nobreak\ \text{e}$ appear in blue.
@@ 3.5.3 Decorate the surface
pair_coeff * * ffield.reax.CHOFe Si O H
@@
The commands above are, once again, similar to the ones of the previous script. Here, the $1\mathrm{e}{-10}$ was added to the denominator of the variable qH to avoid dividing by 0 at the beginning of the simulation, as no hydrogen atoms exists in the simulation domain yet. Finally, let us create a loop with 10 steps, and create two hydrogen atoms at random locations at every step:
@@ 3.6 Tutorial 6: Water adsorption in silica
The objective of this tutorial is to combine molecular dynamics and grand canonical Monte Carlo simulations to compute the adsorption of water molecules in cracked silica material (Fig. 34). This tutorial illustrates the use of the grand canonical ensemble in molecular simulation, an open ensemble where the number of atoms or molecules in the simulation box can vary. By using this combination, we simulate water in a nanoporous SiO2 structure at a specified chemical potential.
@@ 3.6.1 Generation of the silica block
In line with what is done in previous tutorials, the create_atoms commands are used to place 240 Si atoms and 480 O atoms, respectively. This corresponds to an initial density of approximately $2$ g/cm3, which is close to the expected final density of amorphous silica at 300 K.
Now, specify the potential parameters by indicating that the first atom type is Si and the second is O:
@@
Finally, let us implement the annealing procedure which consists of three consecutive runs. This procedure was inspired by Ref. [56]. First, to melt the system, a $10\,\text{ps}$ phase at $T=6000\,\text{K}$ is performed:
@@
n this case, the initial and final target temperatures set for the Nosé-Hoover thermostat is different, causing it to evolve linearly within the number of timesteps evoked in the run command. In the third step, the system is equilibrated at the final desired conditions, $T=300\,\text{K}$ and $p=1\,\text{atm}$, using an anisotropic pressure coupling:
@@
Here, an anisotropic barostat is used. As previously mentioned, anisotropic barostats adjust the dimensions independently, which is generally suitable for a solid phase.
Run the simulation using LAMMPS. From the «Charts» window, the temperature evolution can be observed, showing that it closely follows the desired annealing procedure (Fig. 35 a). The evolution of the box dimensions over time confirms that the box is deforming during the last stage of the simulation (Fig. 35 b). After the simulation completes, the final microstate attained during the dynamics and the system topology will be written to a LAMMPS data file called generate.data which will be located next to generate.lmp (Fig. 36).